OpenStudy (anonymous):

100 = 500 (1 + 0.017/12 ) ^12 t SOLVE FOR T..HOW DO U DO IT

OpenStudy (anonymous):

use logs but i don't know how cause I'm stupid

OpenStudy (anonymous):

@myininaya

jimthompson5910 (jim_thompson5910):

Hopefully I won't get interrupted like last time. That was very very rude. Anyway, 100 = 500(1+0.017/12)^(12t) 100/500 = (1+0.017/12)^(12t) 1/5 = (1+0.017/12)^(12t) ln(1/5) = ln( (1+0.017/12)^(12t) ) ln(1/5) = 12t*ln( 1+0.017/12 ) ln(1/5) = 12t*ln( 1+ 0.00141667) ln(1/5) = 12t*ln( 1.00141667) ln(1/5)/ln( 1.00141667) = 12t -1136.87567652046 = 12t 12t = -1136.87567652046 t = -1136.87567652046/12 t = -94.7396397 So if this was a problem dealing with compound interest, then this means that this answer doesn't make sense. Basically the number on the right MUST be larger than the 500 on the right (since you always make some bit of money) So I would check your numbers again. My guess is that there's a zero missing somewhere.

jimthompson5910 (jim_thompson5910):

Perhaps it should be 1000 = 500(1+0.017/12)^(12t) ?

OpenStudy (anonymous):

Thanks jumbo!! ya i was missing a zero its supposed to be 1000 not 100 lololo

OpenStudy (anonymous):

can't u just use logs or do u have to use ln

jimthompson5910 (jim_thompson5910):

ok, let me fix it

jimthompson5910 (jim_thompson5910):

ln is a log since it's log base 'e'

jimthompson5910 (jim_thompson5910):

do you want me to use normal logs instead?

OpenStudy (anonymous):

wait what is the rules for natural logs i don't get it. its on tomorrows final but i kinda just skimmed it. is it ln base e = power

jimthompson5910 (jim_thompson5910):

\[\Large \ln(x) = \log_{e}(x)\]

jimthompson5910 (jim_thompson5910):

so that shows us that the ln (LN) is a log with base 'e'

OpenStudy (anonymous):